# Multi-fidelity Bayesian Optimization: A Review

### v2

@@ 5.1.4 Multi-step look-ahead acquisition functions
Figure 8 describes BO as a multi-stage decision problem for finding ${\bf x}^{1},{\bf x}^{2},\dots,{\bf x}^{K}$. Our objective is to select ${\bf x}^{1}$ based on the information available in the observed data $\mathcal{D}^{0}$. This selection is influenced by future selections of ${\bf x}^{2},\dots,{\bf x}^{K}$ as we consider the overall outcome of the whole BO process.
@@
where $f^{1}_{\text{H}}=f_{\text{H}}({\bf x}^{1})$ and $\alpha_{1}({\bf x}^{1}|\mathcal{D}^{0})$ is EI given in Eq. 39.
@@
where $\alpha_{K}(\cdot)$ and $\alpha_{K-1}(\cdot)$ represent the $K$- and $(K-1)$-step look-ahead acquisition functions, respectively. The Bellman equation in Eq. 51 indicates that selecting a series of the new design points optimally is to first select ${\bf x}^{1}$ optimally, then select the remaining design points optimally given the objective value from the first selection. Moreover, since $\alpha_{1}({\bf x}^{1}|\mathcal{D}^{0})$ tends to favor exploitation, we can view the first and second terms of Eq. 51 as exploitation and exploration terms, respectively.
@@ 5.2 Acquisition functions considering fidelities
In MF BO, we replace $\hat{f}_{\text{H}}^{k}({\bf x})$ in 9 of Algorithm 1 with one of the MF surrogates described in Section 4. This raises a further question of how to incorporate the information about model fidelities, i.e., $t\in\{1,\dots,T\}$, into the acquisition function so that we can select a new design point and an appropriate computational model for estimating the corresponding objective function value.
@@ 5.2.1 No-fidelity consideration
The acquisition functions for this approach only depend on design variables $\bf x$. Once the new design point ${\bf x}^{k}$ has been found, it is fed to $T$ computational models associated with $T$ fidelities for predictions of the corresponding objective function values. These predictions are used for updating the current solution and MF surrogate. Algorithm 3 shows a pseudo-code for MF BO without considering fidelities. $u(\mathcal{D}^{k})$, $k=0,\dots,K$, in Lines 3 and 10 is defined in Eq. 49. The MF acquisition function in 6 can be any acquisition function of the generic BO. For example, Forrester et al. [16] used EI with $f_{\min}$ selected from the highest fidelity, i.e., $\mathcal{D}^{k}$ in Lines 3 and 10 was fixed at $\mathcal{D}^{K}$. Perdikaris and Karniadakis [131] formulated EI, but with $f_{\min}$ defined as the best objective value among those from all fidelities.
@@
2:$\mathcal{D}^{0}\leftarrow\mathcal{D}_{1}\cup\dots\cup\mathcal{D}_{T}$;
3:$\{{\bf x}_{\min},f_{\min}\}\leftarrow u(\mathcal{D}^{0})$;
@@ 5.2.2 Heuristic approach
The acquisition functions for this approach depend on both $\bf x$ and $t$. They are often derived by using auxiliary functions to modify one of the one-step look-ahead acquisition functions of the generic BO. These auxiliary functions consider the computational cost of each fidelity and/or how the selection of each fidelity for handling the new design point affects the accuracy improvement of MF surrogate. Algorithm 4 shows a pseudo-code for MF BO using the heuristic approach. In Lines 710, a new design point ${\bf x}^{k}_{t}$ is found for each enumerated value of $t$, then the pair $\{{\bf x}^{k}_{t},t\}$ that provides the best acquisition function value ${\alpha}^{k}_{t}=\alpha({\bf x}^{k}_{t},t)$ is selected for the next iteration. Due to its simplicity, the heuristic approach has been widely used in engineering design optimization [see e.g., 15, 201, 137, 144, 156].
@@
2:$\mathcal{D}^{0}\leftarrow\mathcal{D}_{1}\cup\dots\cup\mathcal{D}_{T}$;
3:$\{{\bf x}_{\min},f_{\min}\}\leftarrow u(\mathcal{D}^{0})$;
@@
Attempts to modify Eq. 56 are problem-dependent. One of them is to replace the standard deviation constituting $\text{AEI}({\bf x})$, which depends on ${\bf x}$ only, with a new standard deviation that is a function of both ${\bf x}$ and $t$, given that only two fidelities are considered [70, 142]. Several attempts introduce different forms of $\alpha_{1}({\bf x},t)$. For example, Sacher et al. [149] used $\alpha_{1}({\bf x},t)=\max\left(0,1-\sigma_{f_{\text{H}}}^{2,k+1}({\bf x}|t)/\sigma_{f_{\text{H}}}^{2,k}({\bf x})\right)$, where $\sigma_{f_{\text{H}}}^{2,k+1}({\bf x}|t)$ denotes the predictive variance of the updated MF surrogate, which is formed by adding the new design point generated by fidelity $t$ to the current training dataset. Other attempts define the MF acquisition function as the ratio between an acquisition function of the generic BO and the computational cost per model on the fidelity [160, 154]. Beyond the constraint of most one-step look-ahead heuristic MF acquisition functions, the empirical performance of two-step look-ahead EIs has recently been assessed [137, 157].
@@ 5.2.3 Sequential selection
The sequential selection approach consists of two steps: (1) select the new design point ${\bf x}^{k}$ and (2) select the fidelity $t$ with ${\bf x}^{k}$ found in step (1). The main difference between this approach and the heuristic approach is that the selection of ${\bf x}^{k}$ is independent of that of $t$. This means, the solution improvement is separated from the consideration of computational cost and the accuracy improvement of the MF surrogate. Algorithm 5 shows the pseudo-code for MF BO using the sequential selection approach.
@@
2:$\mathcal{D}^{0}\leftarrow\mathcal{D}_{1}\cup\dots\cup\mathcal{D}_{T}$;
3:$\{{\bf x}_{\min},f_{\min}\}\leftarrow u(\mathcal{D}^{0})$;
@@
[…]
@@
One of the first methods of the sequential approach by Chen et al. [132] used the so-called pre-posterior analysis to formulate a fidelity-query function $\gamma({\bf x}^{k},t)$ (8 of Algorithm 5) after maximizing EI for ${\bf x}^{k}$ in 7. The pre-posterior analysis was to examine how the standard deviation of the MF surrogate at ${\bf x}^{k}$ reduces when fictitious simulation data for each fidelity are added to the currently real training data. Because they used the auto-regressive approach, Chen et al. [132] could generate the samples constituting the fictitious simulation data for fidelity $t$ at ${\bf x}^{k}$ from a Gaussian characterized by the predictive mean and predictive variance at ${\bf x}^{k}$. In this way, a reduction of the standard deviation associated with fidelity $t$ was defined as $\sigma_{f_{\text{H}}}^{k}({\bf x}^{k})-\bar{\sigma}({\bf x}^{k},t)$, where $\bar{\sigma}({\bf x}^{k},t)$ denotes the expected standard deviation from the updated GP posterior after the fictitious simulation data for fidelity $t$ are added to the currently real training data several times. By further considering the computational cost per model on the fidelity, the fidelity-query function was defined as
@@
Alternatively, Kandasamy et al. [117] defined $\gamma({\bf x}^{k},t)$ as the negative value of the computational cost per model on the fidelity. They also imposed two constraints on the maximization of $\gamma({\bf x}^{k},t)$. The first constraint, aiming at promoting exploration, required the posterior variance associated with $({\bf x}^{k},t)$ should be larger than a pre-specified threshold value. The second constraint placed a condition that the maximizer of $\gamma({\bf x}^{k},t)$ should be found in a neighborhood of $T$, which can shrink over time. As a result, the constraints in 9 of Algorithm 5 should consist of the aforementioned two constraints and a bound constraint on $t$ because $t$ is considered a continuous variable [117]; see Section 4.6.1.
In another attempt, Meliani et al. [27] formulated $\gamma({\bf x}^{k},t)$ based on the idea that the use of LF data favors exploration, and that of HF data favors exploitation. Thus, multiple fidelities can be used for updating MF surrogates simultaneously. More specifically, $\gamma({\bf x}^{k},t)$ can be defined as the ratio between the total uncertainty reduction when adding the new data generated from models whose fidelities are not greater than $t$ and the total computational cost of these models. As a result, the models whose fidelities are not greater than $t$ are invoked in 10 of Algorithm 5, which differs from the above-mentioned methods of the sequential approach. This method is a type of the no-fidelity consideration approach when $t=T$.
@@ 5.4 Maximization of acquisition functions
Because it is difficult to examine the convexity of most acquisition functions, we wish to maximize $\alpha({\bf x})$ or $\alpha({\bf x},t)$ using a global optimization algorithm [236]. Nevertheless, robust derivative-based algorithms can be used if an acquisition function has an analytical form. Acquisition functions via Monte-Carlo integration can also be optimized via derivative-based optimization [237]. When maximizing an information-based acquisition function, it may be useful to discretize the design variable space or to use a sampling method. Table 5 lists several optimization algorithms and techniques that are found in the literature to maximize the acquisition functions (and their variants) described in Sections 5.1 and 5.2. Other important aspects of maximizing acquisition functions such as optimizing in latent spaces and on combinatorial domains are discussed in Section 9.2 of Garnett [173].
@@ 6 Illustrative examples
To demonstrate how MF BO can achieve efficient and accurate optimization, we optimize two test functions and the shapes of two standard airfoils using the KOH auto-regressive model combined with EI, PI, or GP-LCB acquisition function (Section 5.1), following the no-fidelity consideration approach (Section 5.2.1). Two fidelity levels are considered for the objective function of each of these problems. The optimization results obtained from MF BO are compared against those from the corresponding generic BO applied to the HF objective function. Note that comparing the optimization results from all the MF BO methods reviewed in this work is impractical because the combination of the MF surrogates and acquisition functions described above results in numerous MF BO algorithms, and there is currently no full-featured MF BO software available for this purpose. Specific comparisons between subsets of MF BO algorithms in various engineering applications are available in the works listed in Table 4.
@@ 6.1 Test functions
Figure 9 shows the medians and interquartile solution ranges from 20 trials of each MF BO and each BO method for the 2d Levy and 6d Hartmann functions. On the 2d Levy function, MF BO with PI achieves the best solutions among the considered methods. The performances of EI and GP-LCB are comparable for both MF BO and BO. On the 6d Hartmann function, all optimization trials from MF BO using EI, PI, and GP-LCB quickly converge at the same solution that is better than those from BO. BO with PI and GP-LCB cannot reach the true optimal solution as they terminate.
@@
$1$ $r_{\text{LE}}$ Leading-edge radius 0.005 0.06
$2$ $x_{\text{up}}$ Upper crest position in horizontal coordinates 0.25 0.5
$3$ $y_{\text{up}}$ Upper crest position in vertical coordinates 0.05 0.15
$4$ $\kappa_{\text{xxup}}$ Upper crest curvature -1 -0.4
$5$ $x_{\text{lo}}$ Lower crest position in horizontal coordinates 0.35 0.5
$6$ $y_{\text{lo}}$ Lower crest position in vertical coordinates -0.12 -0.04
$7$ $\kappa_{\text{xxlo}}$ Lower crest curvature 0.3 1
$8$ $y_{\text{TE}}$ Trailing-edge vertical coordinate -0.02 0.02
$9$ $\Delta y_{\text{TE}}$ Trailing-edge thickness 0 0
@@
$1$ $0.5$ $2$ $6.3\times 10^{6}$ 0.4
$2$ $0.55$ $2.2$ $6.3\times 10^{6}$ 0.2
$3$ $0.5$ $2.5$ $6.3\times 10^{6}$ 0.2
$4$ $0.55$ $2$ $6.3\times 10^{6}$ 0.2
@@ 6.2 Airfoil shapes
Figure 10 shows the parameterization of NACA 4-digit and PARSEC airfoils. The detailed equations are given in Appendix D. The shape of NACA 4-digit airfoils is parameterized by three parameters, namely $c_{\max}$–the height measured from the chord line to the point where the mean camber line has the largest curvature (in absolute value), $x_{\max}$–the position of $c_{\max}$ in horizontal coordinates, and $t_{\max}$–the maximum thickness of the airfoil. Additionally, the cosine spacing is used to generate the coordinates distribution of NACA 4-digit airfoil shapes with open trailing edges. For PARSEC airfoils, the shape is defined by a set of 11 parameters as detailed in Fig. 10(b) and Table 6 [251, 252]. Specific to this work, we fix the trailing-edge thickness at $\Delta y_{\text{TE}}=0$. Thus, there are ten design variables for the shape optimization of PARSEC airfoils. Note that all the length-type parameters are expressed in terms of their ratios to the chord length.
We adopt a multi-point design for the airfoil shapes [78]. This design philosophy aims to achieve airfoil shapes that perform efficiently across various operating conditions, rather than at a single design condition. This design philosophy, therefore, can be considered a type of robust optimization (Section 7.3.1) under uncertainty in operating conditions. Specifically, we consider four distinct operating conditions, as detailed in Table 7, which require the airfoils to operate efficiently at different speeds and angles of attack for a fixed Reynolds number. As a result, we define the HF objective function for the shape optimization problem as the weighted sum of negative lift-to-drag ratios across the four operating conditions. The optimization problem for NACA 4-digit and PARSEC airfoils is formulated as follows:
@@
[…]
Here ${\bf x}=\left[c_{\max},x_{\max},t_{\max}\right]^{\intercal}$ and ${\bf x}=\left[r_{\text{LE}},x_{\text{up}},y_{\text{up}},\kappa_{\text{xxup}},x_{\text{lo}},y_{\text{lo}},\kappa_{\text{xxlo}},y_{\text{TE}},\alpha_{\text{TE}},\beta_{\text{TE}}\right]^{\intercal}$ are the vectors of design variables for NACA 4-digit and PARSEC airfoils, respectively. $\underline{{\bf x}}$ and $\overline{{\bf x}}$ are the lower and upper bounds of ${\bf x}$, where $\underline{{\bf x}}=[0,0,0.1]^{\intercal}$ and $\overline{{\bf x}}=[0.08,0.8,0.25]^{\intercal}$ for NACA 4-digit airfoils, while those for PARSEC airfoils are given in Table 6. The HF objective function reads $f_{\text{H}}({\bf x})=-\sum_{i=1}^{4}w_{i}c_{\text{L},i}({\bf x})/c_{\text{D},i}({\bf x})$, where $c_{\text{L},i}({\bf x})$ and $c_{\text{D},i}({\bf x})$ represent the lift and drag coefficients at the $i$-th operating condition, respectively, $w_{i}$ listed in Table 7 is the weight for the $i$-th operating condition, and the negative sign is to transform a maximization problem of the lift-to-drag ratio into a minimization problem of $f_{\text{H}}({\bf x})$.
@@
For optimizing NACA 4-digit airfoils, we generate ten initial datasets for ten different trials of each optimization method. Each dataset has 15 initial HF $+$ 15 initial LF observations for MF BO, and 20 HF observations for BO. We also generate ten initial datasets for optimizing PARSEC airfoils, each for MF BO having 20 initial HF $+$ 20 initial LF observations, and 30 HF observations for BO. Note that there is no strict guideline for determining the number of initial HF observations to start BO and MF BO. We can start each method with an empty dataset [173], a dataset of $10d$ HF observations [168, 255], or a dataset of any size between these two extremes, where $d$ is the number of design variables. Here we initiate MF BO with fewer HF observations than BO as we hypothesize that once both methods reach a sufficiently large number of HF observations, MF BO, leveraging useful information from LF observations at a minor extra cost, would yield a solution that is comparable to (when both methods converge) or better than that of BO, despite requiring fewer initial HF observations. This is justified because we adopt the no-fidelity consideration MF BO approach that calls both HF and LF models after each optimization iteration. In each optimization iteration, we record the best-found observation value $f_{\min}$ of the HF objective function and the corresponding design variables.
@@
Figure 11 shows the medians and interquartile ranges of solutions from ten trials of each MF BO and each BO method for NACA 4-digit and PARSEC airfoils. As expected, MF BO outperforms BO in both solution quality and the consistency of the final solutions. For NACA 4-digit airfoils, the three considered MF BO methods are comparable. For PARSEC 4-digit airfoils, MF BO with GP-LCB yields the best solution. This variation in the performance of the MF BO methods across the two optimization problems highlights that no acquisition function works well on all classes of problems; see Section 5.3.
@@
Figures 12 and 13 illustrate the airfoil shapes and corresponding pressure coefficient distributions, $C_{\text{p}}$, for the best designs of NACA 4-digit and PARSEC airfoils obtained from each optimization method. We see that the trial of each MF BO method that provides the best design considerably changes the airfoil shape from an inefficient design to an optimized design with an improved lift characteristic under four operating conditions. While the best NACA 4-digit airfoil designs by MF BO and BO with the same acquisition function are almost identical, the best PARSEC airfoil designs by these methods exhibit slight differences in their geometries.

### v3

@@ 5.1.4 Multi-step look-ahead acquisition functions
Figure˜8 describes BO as a multi-stage decision problem for finding ${\bf x}^{1},{\bf x}^{2},\dots,{\bf x}^{K}$. Our objective is to select ${\bf x}^{1}$ based on the information available in the observed data $\mathcal{D}^{0}$. This selection is influenced by future selections of ${\bf x}^{2},\dots,{\bf x}^{K}$ as we consider the overall outcome of the whole BO process.
@@
where $f^{1}_{\text{H}}=f_{\text{H}}({\bf x}^{1})$ and $\alpha_{1}({\bf x}^{1}|\mathcal{D}^{0})$ is EI given in Eq.˜39.
@@
where $\alpha_{K}(\cdot)$ and $\alpha_{K-1}(\cdot)$ represent the $K$- and (K1)(K-1)-step look-ahead acquisition functions, respectively. The Bellman equation in Eq.˜51 indicates that selecting a series of the new design points optimally is to first select ${\bf x}^{1}$ optimally, then select the remaining design points optimally given the objective value from the first selection. Moreover, since $\alpha_{1}({\bf x}^{1}|\mathcal{D}^{0})$ tends to favor exploitation, we can view the first and second terms of Eq.˜51 as exploitation and exploration terms, respectively.
@@ 5.2 Acquisition functions considering fidelities
In MF BO, we replace $\hat{f}_{\text{H}}^{k}({\bf x})$ in Line 9 of Algorithm˜1 with one of the MF surrogates described in Section˜4. This raises a further question of how to incorporate the information about model fidelities, i.e., $t\in\{1,\dots,T\}$, into the acquisition function so that we can select a new design point and an appropriate computational model for estimating the corresponding objective function value.
@@ 5.2.1 No-fidelity consideration
The acquisition functions for this approach only depend on design variables $\bf x$. Once the new design point ${\bf x}^{k}$ has been found, it is fed to $T$ computational models associated with $T$ fidelities for predictions of the corresponding objective function values. These predictions are used for updating the current solution and MF surrogate. Algorithm˜3 shows a pseudo-code for MF BO without considering fidelities. $u(\mathcal{D}^{k})$, $k=0,\dots,K$, in Lines 3 and 10 is defined in Eq.˜49. The MF acquisition function in Line 6 can be any acquisition function of the generic BO. For example, Forrester et al. [16] used EI with $f_{\min}$ selected from the highest fidelity, i.e., $\mathcal{D}^{k}$ in Lines 3 and 10 was fixed at $\mathcal{D}^{K}$. Perdikaris and Karniadakis [131] formulated EI, but with $f_{\min}$ defined as the best objective value among those from all fidelities.
@@
[…]
@@ 5.2.2 Heuristic approach
The acquisition functions for this approach depend on both $\bf x$ and $t$. They are often derived by using auxiliary functions to modify one of the one-step look-ahead acquisition functions of the generic BO. These auxiliary functions consider the computational cost of each fidelity and/or how the selection of each fidelity for handling the new design point affects the accuracy improvement of MF surrogate. Algorithm˜4 shows a pseudo-code for MF BO using the heuristic approach. In Lines 710, a new design point ${\bf x}^{k}_{t}$ is found for each enumerated value of $t$, then the pair $\{{\bf x}^{k}_{t},t\}$ that provides the best acquisition function value ${\alpha}^{k}_{t}=\alpha({\bf x}^{k}_{t},t)$ is selected for the next iteration. Due to its simplicity, the heuristic approach has been widely used in engineering design optimization [see e.g., 15, 201, 137, 144, 156].
@@
[…]
@@
Attempts to modify Eq.˜56 are problem-dependent. One of them is to replace the standard deviation constituting $\text{AEI}({\bf x})$, which depends on $\bf x$} only, with a new standard deviation that is a function of both $\bf x$} and $t$, given that only two fidelities are considered [70, 142]. Several attempts introduce different forms of $\alpha_{1}({\bf x},t)$. For example, Sacher et al. [149] used $\alpha_{1}({\bf x},t)=\max\left(0,1-\sigma_{f_{\text{H}}}^{2,k+1}({\bf x}|t)/\sigma_{f_{\text{H}}}^{2,k}({\bf x})\right)$, where $\sigma_{f_{\text{H}}}^{2,k+1}({\bf x}|t)$ denotes the predictive variance of the updated MF surrogate, which is formed by adding the new design point generated by fidelity $t$ to the current training dataset. Other attempts define the MF acquisition function as the ratio between an acquisition function of the generic BO and the computational cost per model on the fidelity [160, 154]. Beyond the constraint of most one-step look-ahead heuristic MF acquisition functions, the empirical performance of two-step look-ahead EIs has recently been assessed [137, 157].
@@ 5.2.3 Sequential selection
The sequential selection approach consists of two steps: (1) select the new design point ${\bf x}^{k}$ and (2) select the fidelity $t$ with ${\bf x}^{k}$ found in step (1). The main difference between this approach and the heuristic approach is that the selection of ${\bf x}^{k}$ is independent of that of $t$. This means, the solution improvement is separated from the consideration of computational cost and the accuracy improvement of the MF surrogate. Algorithm˜5 shows the pseudo-code for MF BO using the sequential selection approach.
@@
[…]
@@
9:  $t\leftarrow\underset{t}{\mathrm{argmax}}\ \ \gamma({\bf x}^{k},t)$ s.t. $t\in\{1,\dots,T\}$;
@@
One of the first methods of the sequential approach by Chen et al. [132] used the so-called pre-posterior analysis to formulate a fidelity-query function $\gamma({\bf x}^{k},t)$ (Line 8 of Algorithm˜5) after maximizing EI for ${\bf x}^{k}$ in Line 7. The pre-posterior analysis was to examine how the standard deviation of the MF surrogate at ${\bf x}^{k}$ reduces when fictitious simulation data for each fidelity are added to the currently real training data. Because they used the auto-regressive approach, Chen et al. [132] could generate the samples constituting the fictitious simulation data for fidelity $t$ at ${\bf x}^{k}$ from a Gaussian characterized by the predictive mean and predictive variance at ${\bf x}^{k}$. In this way, a reduction of the standard deviation associated with fidelity $t$ was defined as $\sigma_{f_{\text{H}}}^{k}({\bf x}^{k})-\bar{\sigma}({\bf x}^{k},t)$, where $\bar{\sigma}({\bf x}^{k},t)$ denotes the expected standard deviation from the updated GP posterior after the fictitious simulation data for fidelity $t$ are added to the currently real training data several times. By further considering the computational cost per model on the fidelity, the fidelity-query function was defined as
@@
Alternatively, Kandasamy et al. [117] defined $\gamma({\bf x}^{k},t)$ as the negative value of the computational cost per model on the fidelity. They also imposed two constraints on the maximization of $\gamma({\bf x}^{k},t)$. The first constraint, aiming at promoting exploration, required the posterior variance associated with $({\bf x}^{k},t)$ should be larger than a pre-specified threshold value. The second constraint placed a condition that the maximizer of $\gamma({\bf x}^{k},t)$ should be found in a neighborhood of $T$, which can shrink over time. As a result, the constraints in Line 9 of Algorithm˜5 should consist of the aforementioned two constraints and a bound constraint on $t$ because $t$ is considered a continuous variable [117]; see Section˜4.6.1.
In another attempt, Meliani et al. [27] formulated $\gamma({\bf x}^{k},t)$ based on the idea that the use of LF data favors exploration, and that of HF data favors exploitation. Thus, multiple fidelities can be used for updating MF surrogates simultaneously. More specifically, $\gamma({\bf x}^{k},t)$ can be defined as the ratio between the total uncertainty reduction when adding the new data generated from models whose fidelities are not greater than $t$ and the total computational cost of these models. As a result, the models whose fidelities are not greater than $t$ are invoked in Line 10 of Algorithm˜5, which differs from the above-mentioned methods of the sequential approach. This method is a type of the no-fidelity consideration approach when $t=T$.
@@ 5.4 Maximization of acquisition functions
Because it is difficult to examine the convexity of most acquisition functions, we wish to maximize $\alpha({\bf x})$ or $\alpha({\bf x},t)$ using a global optimization algorithm [236]. Nevertheless, robust derivative-based algorithms can be used if an acquisition function has an analytical form. Acquisition functions via Monte-Carlo integration can also be optimized via derivative-based optimization [237]. When maximizing an information-based acquisition function, it may be useful to discretize the design variable space or to use a sampling method. Table˜5 lists several optimization algorithms and techniques that are found in the literature to maximize the acquisition functions (and their variants) described in Sections 5.1 and 5.2. Other important aspects of maximizing acquisition functions such as optimizing in latent spaces and on combinatorial domains are discussed in Section 9.2 of Garnett [173].
@@ 6 Illustrative examples
To demonstrate how MF BO can achieve efficient and accurate optimization, we optimize two test functions and the shapes of two standard airfoils using the KOH auto-regressive model combined with EI, PI, or GP-LCB acquisition function (Section˜5.1), following the no-fidelity consideration approach (Section˜5.2.1). Two fidelity levels are considered for the objective function of each of these problems. The optimization results obtained from MF BO are compared against those from the corresponding generic BO applied to the HF objective function. Note that comparing the optimization results from all the MF BO methods reviewed in this work is impractical because the combination of the MF surrogates and acquisition functions described above results in numerous MF BO algorithms, and there is currently no full-featured MF BO software available for this purpose. Specific comparisons between subsets of MF BO algorithms in various engineering applications are available in the works listed in Table˜4.
@@ 6.1 Test functions
Figure˜9 shows the medians and interquartile solution ranges from 20 trials of each MF BO and each BO method for the 2d Levy and 6d Hartmann functions. On the 2d Levy function, MF BO with PI achieves the best solutions among the considered methods. The performances of EI and GP-LCB are comparable for both MF BO and BO. On the 6d Hartmann function, all optimization trials from MF BO using EI, PI, and GP-LCB quickly converge at the same solution that is better than those from BO. BO with PI and GP-LCB cannot reach the true optimal solution as they terminate.
@@
11 $r_{\text{LE}}$ Leading-edge radius 0.005 0.06
22 $x_{\text{up}}$ Upper crest position in horizontal coordinates 0.25 0.5
33 $y_{\text{up}}$ Upper crest position in vertical coordinates 0.05 0.15
44 $\kappa_{\text{xxup}}$ Upper crest curvature -1 -0.4
55 $x_{\text{lo}}$ Lower crest position in horizontal coordinates 0.35 0.5
66 $y_{\text{lo}}$ Lower crest position in vertical coordinates -0.12 -0.04
77 $\kappa_{\text{xxlo}}$ Lower crest curvature 0.3 1
88 $y_{\text{TE}}$ Trailing-edge vertical coordinate -0.02 0.02
99 $\Delta y_{\text{TE}}$ Trailing-edge thickness 0 0
@@
11 $0.5$ 22 $6.3\times 10^{6}$ 0.4
22 $0.55$ $2.2$ $6.3\times 10^{6}$ 0.2
33 $0.5$ $2.5$ $6.3\times 10^{6}$ 0.2
44 $0.55$ 22 $6.3\times 10^{6}$ 0.2
@@ 6.2 Airfoil shapes
Figure˜10 shows the parameterization of NACA 4-digit and PARSEC airfoils. The detailed equations are given in Appendix D. The shape of NACA 4-digit airfoils is parameterized by three parameters, namely $c_{\max}$–the height measured from the chord line to the point where the mean camber line has the largest curvature (in absolute value), $x_{\max}$–the position of $c_{\max}$ in horizontal coordinates, and $t_{\max}$–the maximum thickness of the airfoil. Additionally, the cosine spacing is used to generate the coordinates distribution of NACA 4-digit airfoil shapes with open trailing edges. For PARSEC airfoils, the shape is defined by a set of 11 parameters as detailed in Fig.˜10(b) and Table˜6 [251, 252]. Specific to this work, we fix the trailing-edge thickness at $\Delta y_{\text{TE}}=0$. Thus, there are ten design variables for the shape optimization of PARSEC airfoils. Note that all the length-type parameters are expressed in terms of their ratios to the chord length.
We adopt a multi-point design for the airfoil shapes [78]. This design philosophy aims to achieve airfoil shapes that perform efficiently across various operating conditions, rather than at a single design condition. This design philosophy, therefore, can be considered a type of robust optimization (Section˜7.3.1) under uncertainty in operating conditions. Specifically, we consider four distinct operating conditions, as detailed in Table˜7, which require the airfoils to operate efficiently at different speeds and angles of attack for a fixed Reynolds number. As a result, we define the HF objective function for the shape optimization problem as the weighted sum of negative lift-to-drag ratios across the four operating conditions. The optimization problem for NACA 4-digit and PARSEC airfoils is formulated as follows:
@@
s.t. ${\bf x}\in\left[\underline{{\bf x}},\overline{{\bf x}}\right].$
Here ${\bf x}=\left[c_{\max},x_{\max},t_{\max}\right]^{\intercal}$ and ${\bf x}=\left[r_{\text{LE}},x_{\text{up}},y_{\text{up}},\kappa_{\text{xxup}},x_{\text{lo}},y_{\text{lo}},\kappa_{\text{xxlo}},y_{\text{TE}},\alpha_{\text{TE}},\beta_{\text{TE}}\right]^{\intercal}$ are the vectors of design variables for NACA 4-digit and PARSEC airfoils, respectively. $\underline{{\bf x}}$ and $\overline{{\bf x}}$ are the lower and upper bounds of $\bf x$}, where $\underline{{\bf x}}=[0,0,0.1]^{\intercal}$ and $\overline{{\bf x}}=[0.08,0.8,0.25]^{\intercal}$ for NACA 4-digit airfoils, while those for PARSEC airfoils are given in Table˜6. The HF objective function reads $f_{\text{H}}({\bf x})=-\sum_{i=1}^{4}w_{i}c_{\text{L},i}({\bf x})/c_{\text{D},i}({\bf x})$, where $c_{\text{L},i}({\bf x})$ and $c_{\text{D},i}({\bf x})$ represent the lift and drag coefficients at the $i$-th operating condition, respectively, $w_{i}$ listed in Table˜7 is the weight for the $i$-th operating condition, and the negative sign is to transform a maximization problem of the lift-to-drag ratio into a minimization problem of $f_{\text{H}}({\bf x})$.
@@
For optimizing NACA 4-digit airfoils, we generate ten initial datasets for ten different trials of each optimization method. Each dataset has 15 initial HF ++ 15 initial LF observations for MF BO, and 20 HF observations for BO. We also generate ten initial datasets for optimizing PARSEC airfoils, each for MF BO having 20 initial HF ++ 20 initial LF observations, and 30 HF observations for BO. Note that there is no strict guideline for determining the number of initial HF observations to start BO and MF BO. We can start each method with an empty dataset [173], a dataset of $10d$ HF observations [168, 255], or a dataset of any size between these two extremes, where $d$ is the number of design variables. Here we initiate MF BO with fewer HF observations than BO as we hypothesize that once both methods reach a sufficiently large number of HF observations, MF BO, leveraging useful information from LF observations at a minor extra cost, would yield a solution that is comparable to (when both methods converge) or better than that of BO, despite requiring fewer initial HF observations. This is justified because we adopt the no-fidelity consideration MF BO approach that calls both HF and LF models after each optimization iteration. In each optimization iteration, we record the best-found observation value $f_{\min}$ of the HF objective function and the corresponding design variables.
@@
Figure˜11 shows the medians and interquartile ranges of solutions from ten trials of each MF BO and each BO method for NACA 4-digit and PARSEC airfoils. As expected, MF BO outperforms BO in both solution quality and the consistency of the final solutions. For NACA 4-digit airfoils, the three considered MF BO methods are comparable. For PARSEC 4-digit airfoils, MF BO with GP-LCB yields the best solution. This variation in the performance of the MF BO methods across the two optimization problems highlights that no acquisition function works well on all classes of problems; see Section˜5.3.
@@
Figures˜12 and 13 illustrate the airfoil shapes and corresponding pressure coefficient distributions, $C_{\text{p}}$, for the best designs of NACA 4-digit and PARSEC airfoils obtained from each optimization method. We see that the trial of each MF BO method that provides the best design considerably changes the airfoil shape from an inefficient design to an optimized design with an improved lift characteristic under four operating conditions. While the best NACA 4-digit airfoil designs by MF BO and BO with the same acquisition function are almost identical, the best PARSEC airfoil designs by these methods exhibit slight differences in their geometries.